# A Transfer Learning-Based Method for Water Body Segmentation in Remote Sensing Imagery: A Case Study of the Zhada Tulin Area

### v1

@@ 3.2 Comparative Analysis of Transfer Learning
Table 2: Performance comparison of the models on the validation set of target dataset B.
Model Class IoU (%) F1-score (%) Precision (%) Recall (%)
A1 (Direct Transfer) Background 72.06 83.76 75.25 94.43
Water 25.50 40.64 69.24 28.76
B-scratch (SegFormer) Background 69.58 82.06 77.38 87.35
Water 37.47 54.51 64.40 47.26
B-scratch (U-Net) Background 72.00 84.34 83.11 84.34
Water 48.82 65.61 66.65 64.61
A2 (Fine-tuned) Background 77.59 87.38 89.80 85.09
Water 64.84 78.67 75.25 82.42
@@
The updated results, as shown in Table 2, reveal performance differences across strategies. The direct transfer model (A1) performed the poorest, with a water IoU of 25.50% and a recall of 28.76%. This drop (compared to its 68.80% IoU in the source domain) highlights the domain shift issue.
@@ 4 Discussion
Figure 6 compares the performance of three strategies in a typical Zhada Tulin scene, where the spectral features between water bodies and the background (Fig. 6a) are highly similar.
The result from directly applying the source domain model A1 (Fig. 6c) confirms the domain shift issue. The prediction consists of fragmented and disordered patches, losing the river’s overall morphology. This indicates that features learned by Model A1 in the source domain do not generalize well to the new scene.
The result from the SegFormer model trained from scratch (B-scratch (Seg)) (Fig. 6d) highlights the challenges of small-sample learning. With only 162 training samples, the model fails to learn robust features for distinguishing water from background, often misclassifying local non-essential elements (e.g., mountain shadows or sediment textures) as water. This leads to incomplete convergence and poor performance on both training and validation sets. For comparison, the U-Net model trained from scratch (B-scratch (U-Net))[28] achieved a water IoU of 48.82%, outperforming the SegFormer scratch result of 37.47%.
In contrast, the proposed transfer learning strategy (Model A2) produces better results (Fig. 6e). The predictions align closely with the ground truth (Fig. 6b), preserving the river’s structure, connectivity, and minimizing background misclassifications.
The experiments reveal a notable pattern: training the SegFormer model with ImageNet pre-trained weights on the target domain (B-scratch (Seg)) yields an IoU of 37.47%, lower than the 48.82% from the simpler U-Net trained from random initialization. This can be attributed to the small sample size in the Zhada dataset, where the complex Transformer architecture of SegFormer struggles with limited data, resulting in suboptimal fitting. The U-Net, with its simpler CNN structure, shows greater robustness under such constraints.
However, both baselines underperform compared to the proposed transfer learning model A2 (IoU=64.84%). This underscores that model architecture alone—whether SegFormer or U-Net—cannot address the core issues. The key lies in the two-stage paradigm, which transfers domain-specific knowledge from dataset A to enhance the SegFormer model, overcoming small-sample and domain shift limitations. This approach offers a straightforward alternative to more intricate domain adaptation techniques, requiring fewer unlabeled samples.
The study’s outcomes stem from the transfer learning strategy’s role in knowledge reconstruction [25, 40]. Stage 1 pre-training equips Model A1 with general low-level features (e.g., color, texture, edges) from diverse source data. Direct application to Zhada Tulin, however, falls short due to unique local characteristics. Stage 2 fine-tuning adapts high-level semantics using targeted target-domain samples, enabling the model to handle specific distributions like sediment-influenced spectra and complex backgrounds. Radiation differences between GF-7 and GF-2 data were retained to mimic real sensor variations, improving method robustness.
It is important to note that the source domain imagery was not pristine; many images contained significant masked-out regions, typically corresponding to cloud cover, cloud shadows, or sensor artifacts. These areas, visually distinct and containing no valid data values, increase the complexity of the training environment. The model must not only learn to identify water bodies but also to effectively ignore these large, irregular "no-data" zones, which adds to the challenge of learning robust and generalizable features and makes the pre-trained model more resilient to real-world data imperfections.
[…]
Practical Implications and Limitations Beyond its technical contributions, the method shows promise in real-world remote sensing contexts. It enables cost-effective extraction of water body boundaries in remote plateau areas. For instance, in ecological monitoring, it can be used to track seasonal and long-term variations in water areas to inform environmental protection policies. In disaster risk management, it aids in detecting dynamic changes in glacial lake shapes, contributing to early warning systems. For water resource evaluation, it facilitates quantitative analysis of regional water volumes and distributions, proving especially valuable in resource-scarce yet high-demand remote regions. These aspects underscore the strategy’s adaptability, particularly for plateau-specific challenges like high sediment loads and intricate background interferences.
Although the proposed method improves performance, limitations persist. The water IoU of 64.84% for Model A2, while a significant improvement, suggests room for enhancement, particularly in highly interfering backgrounds. This is partly attributable to the small size of dataset B, despite data augmentation. Relative to benchmarks in similar studies (e.g.,  70% IoU on the LoveDA dataset [35]), the results are reasonable for a challenging small-sample plateau setting, but additional data or more advanced augmentation could be beneficial. This study focuses on the pre-train-fine-tune paradigm; future work could incorporate advanced domain adaptation methods [8, 33, 19, 9, 30, 12, 27], such as adversarial or self-supervised approaches, to further reduce the domain gap with minimal labeling. Extending the framework to other remote sensing tasks, such as building or vegetation extraction, is also a promising direction.
@@ 5 Conclusion
This study develops a two-stage transfer learning strategy based on the SegFormer model to mitigate domain shift and small-sample challenges in remote sensing image water body segmentation. Validated through the challenging Zhada Tulin case, the proposed strategy elevates the water IoU from 25.50% in direct transfer to 64.84%, substantially surpassing scratch-trained baselines, including SegFormer (37.47%) and the simpler U-Net (48.82%). These findings highlight the effectiveness of pre-training on a diverse source domain followed by targeted fine-tuning. This approach successfully transfers general water features while adapting to unique local spectral and geomorphological distributions, thereby achieving reliable segmentation under limited-data conditions. In conclusion, this work demonstrates the practical benefits of knowledge reconstruction for remote sensing and offers a robust, cost-effective, and replicable framework highly valuable for thematic information extraction in similar data-scarce environments.
Declarations
[…]
Author Contributions All authors contributed to the study conception and design. Haonan Chen was responsible for the initial investigation, data acquisition, and preprocessing. Building on this foundational work, the formal analysis, methodology development, software implementation, and validation experiments were carried out by Xin Tong. Both authors contributed to the writing, review, and editing of the manuscript and have approved the final version.
[…]
Funding The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
[…]
Data Availability The datasets generated and/or analysed during the current study are available from the first author (H.C.) on reasonable request.
[…]
Code Availability The source code developed for this study is available from the corresponding author (X.T.) on reasonable request.
[…]
Conflict of Interest The authors declare that they have no competing interests.
[…]
Ethics Approval Not applicable.

### v2

@@ 3.2 Comparative Analysis of Transfer Learning
Performance comparison of the models on the validation set of target dataset B. Model Class IoU (%) F1-score (%) Precision (%) Recall (%) A1 (Direct Transfer) Background 72.06 83.76 75.25 94.43 Water 25.50 40.64 69.24 28.76 B-scratch (SegFormer) Background 69.58 82.06 77.38 87.35 Water 37.47 54.51 64.40 47.26 B-scratch (U-Net) Background 72.00 84.34 83.11 84.34 Water 48.82 65.61 66.65 64.61 A2 (Fine-tuned) Background 77.59 87.38 89.80 85.09 Water 64.84 78.67 75.25 82.42
The updated results, as shown in Table 3.2, reveal performance differences across strategies. The direct transfer model (A1) performed the poorest, with a water IoU of 25.50% and a recall of 28.76%. This drop (compared to its 68.80% IoU in the source domain) highlights the domain shift issue.
@@ 4 Discussion
Our experimental results highlight the critical role of transfer learning in overcoming the dual challenges of domain shift and data scarcity in remote sensing segmentation. The direct transfer model (A1) and the scratch-trained models (B-scratch SegFormer and U-Net) all yielded suboptimal performance, confirming that neither a general model nor a complex architecture alone is sufficient. The proposed two-stage fine-tuning strategy, however, achieved a water IoU of 64.84%, demonstrating its effectiveness. The visual results in Figure 6 corroborate these findings, showing that only the fine-tuned model (A2) could accurately restore the river’s morphology against a complex background.
The superior performance of our approach stems from its effective knowledge reconstruction process (Pan and Yang 2010; Yosinski et al. 2014). The source-domain pre-training endowed the model with robust low-level feature extraction capabilities, honed on diverse data that even included significant artifacts like cloud masks. This pre-training made the model resilient to real-world data imperfections. Subsequently, the target-domain fine-tuning allowed the model to adapt its high-level semantic understanding to the specific spectral and geomorphological characteristics of the Zhada Tulin, including high turbidity and the unique sensor properties of GF-2 imagery. This ”general-to-specific” adaptation proved crucial for success.
4.1 Geoscientific Implications of the Segmentation Results
Beyond the technical advancements, the high-precision water body map generated by our model provides significant insights into the geoscientific characteristics of the Zhada Tulin area (Li et al. 2013; Zhu et al. 2019; Chen et al. 2009). The segmentation results reveal that the water systems exhibit a distinct dual control by tectonics and lithology. The main channel of the Xiangquan River follows a NW-SE trend, strictly controlled by regional fault structures, while its tributaries are primarily incised into softer lacustrine sediments, forming deep canyons. This drainage pattern serves as a microcosm of the coupled processes of tectonic uplift and fluvial incision during the Quaternary, reflecting the geomorphological evolution of the Tibetan Plateau.
More importantly, our high-precision extraction reveals the highly concentrated nature of water resources in this arid region. Statistical analysis indicates that over 80% of the water surface area is confined to less than 20% of the total river channel length, mainly along the Xiangquan River’s main stem and its three primary tributaries. This extremely uneven distribution creates a pronounced ’corridor effect’, where the river valleys form relatively humid oases that support the region’s limited vegetation cover and wildlife habitats, while the vast surrounding earth forest plateaus remain largely barren. From a climate resilience and disaster risk management perspective, this finding underscores the high dependency of local human activities, such as seasonal grazing, on these few critical water sources. The concentrated water distribution pattern also has important implications for flood risk assessment and early warning systems, as flash floods during intense precipitation events are likely to be concentrated in these narrow corridor zones. Therefore, establishing a long-term monitoring system based on our remote sensing approach is of great practical significance for adapting to climate change and maintaining ecological balance in the region,while also supporting disaster preparedness efforts in this climate-sensitive area.
4.2 Limitations and Future Directions
Despite the promising results, this study has limitations. The water IoU of 64.84%, while a significant improvement, indicates that challenges remain, particularly in scenes with extreme background interference. This is partly attributable to the limited size of the target domain dataset. Although our results are reasonable for a challenging small-sample plateau setting when compared to benchmarks on other datasets (e.g.,  70% IoU on LoveDA (Wang et al. 2021)), further improvements could be achieved with more data.
Our study primarily validated the classic pre-train-fine-tune paradigm. Future research should explore two key directions. First, developing a multi-temporal monitoring system to analyze long-term water dynamics, which would provide deeper insights into the hydrological response to climate change and support enhanced early warning capabilities for climate-related hazards.Second, incorporating more advanced domain adaptation techniques (Ganin and Lempitsky 2015; Tzeng et al. 2017; Long et al. 2015; Ghifary et al. 2016; Saito et al. 2018; Hoffman et al. 2018; Pei et al. 2018), such as adversarial or self-supervised learning, could further bridge the domain gap with even less labeled data. Extending this framework to other remote sensing tasks, like building or vegetation extraction, also presents valuable opportunities for advancing Earth observation capabilities that support sustainability and climate resilience applications.
@@ 5 Conclusion
This study proposed and validated a two-stage transfer learning strategy based on the SegFormer model to address the critical challenges of domain shift and small-sample learning in remote sensing water body segmentation. By pre-training on a diverse source domain and subsequently fine-tuning on a specific, challenging target domain—the Zhada Tulin—our approach demonstrated significant performance gains. The water body segmentation IoU on the target validation set surged from 25.50% (for direct transfer) to 64.84%, substantially outperforming scratch-trained baselines.
Beyond the technical achievement, this research underscores the geoscientific value of high-precision segmentation. The generated water map revealed the highly concentrated spatial distribution of water resources in the Zhada Tulin, a key insight for understanding the hydrological processes and delicate ecological balance in this arid plateau region. This work not only provides a robust and replicable framework for thematic information extraction in data-scarce environments but also offers vital technical support for monitoring the headwaters of transboundary rivers and assessing water security under a changing climate. In conclusion, our study effectively bridges advanced deep learning techniques with pressing geoscientific applications, offering a valuable paradigm for future remote sensing research.
Author Contributions
All authors contributed to the study conception and design. Haonan Chen was responsible for the initial investigation, data acquisition, and preprocessing. Building on this foundational work, the formal analysis, methodology development, software implementation, and validation experiments were carried out by Xin Tong. Both authors contributed to the writing, review, and editing of the manuscript and have approved the final version.
Disclosure of interest
The authors report no conflict of interest.
Data availability statement
The datasets generated and/or analysed during the current study are available from the first author (H.C.) on reasonable request.
Code availability
The source code developed for this study is available from the corresponding author (X.T.) on reasonable request.